# On the Number of Control Nodes in Boolean Networks with Degree Constraints

### v1

@@ 5 Boolean networks consisting of AND/OR functions and negations
More generally, according to Theorem 8, for each $l=l^{*}$, $\frac{k-(l^{*}+1)}{2k-(l^{*}+1)}n+\frac{1}{2k-(l^{*}+1)}\left(\sum_{j=1}^{(l^{*}+1)-2}j\cdot|M_{j}|\right)+\frac{(l^{*}+1)-1}{2k-(l^{*}+1)}\left(\sum_{j=(l^{*}+1)-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)>\frac{k-l^{*}}{2k-l^{*}}n+\frac{1}{2k-l^{*}}\left(\sum_{j=1}^{l^{*}-2}j\cdot|M_{j}|\right)+\frac{l^{*}-1}{2k-l^{*}}\left(\sum_{j=l^{*}-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)$, when $\sum_{j=1}^{l^{*}-1}j\cdot|M_{j}|+(2k-1)(|M_{l^{*}}|+\cdots+|M_{\lfloor\frac{k}{2}\rfloor}|)>kn$. Similarly, it can be seen that the larger $|M_{j}|,~{}j=1,\ldots,\lfloor\frac{k}{2}\rfloor$, the larger the general lower bound on the size of control node set, and the influence of the values of $|M_{1}|,\ldots,|M_{\lfloor\frac{k}{2}\rfloor}|$ on the value of the general lower bound is from small to large, that is, $|M_{\lfloor\frac{k}{2}\rfloor}|$ has the greatest impact on the value of the general lower bound.
@@ 6 Boolean networks consisting of nested canalyzing functions
Finally, we consider the case where the nested canalyzing function is assigned to each node, and we refer to this BN as a $k$-$k$-NC-BN.
@@
When $(n\mod k)=0$, we partition $[x_{1},x_{2},\ldots,x_{n}]$ into $\frac{n}{k}$ blocks:
$[x_{1},x_{2},\ldots,x_{k}],[x_{k+1},x_{k+2},\ldots,x_{2k}],\ldots,$
@@ 7 Conclusion
  • According to Theorem 2, when $t^{*}=2$, the general lower bound on the size of the control node set is $\frac{n}{2}$. On the other hand, the best case upper bound for 2-2-XOR-BNs is $\frac{n}{2}$ mentioned in Proposition 2, which means that the general lower bound$=$best case upper bound in this case, i.e., these bounds are optimal.
@@
  • More generally, according to Remark 10, it can be seen that the larger of the values of $|M_{j}|,~{}j=1,\ldots,\lfloor\frac{k}{2}\rfloor$, the larger the general lower bound given in Theorem 8, and the influence of the values of $|M_{1}|,\ldots,|M_{\lfloor\frac{k}{2}\rfloor}|$ on the value of the general lower bound is from small to large, that is, $|M_{\lfloor\frac{k}{2}\rfloor}|$ has the greatest impact on the value of the general lower bound.
@@
However, there are still some open issues that need to be resolved in the future. First, the gap between the bounds given by Proposition 2 and Theorem 3 is significant, so it is necessary to narrow this gap further. Moreover, a general upper bound or a worst case lower bound for $k$-$k$-AND-BN with negation (including the case of $k=2$) is not given in the paper, so it is necessary to derive these bounds further.

### v2

@@ 5 Boolean networks consisting of AND/OR functions and negations
More generally, according to Theorem 8, the number of control nodes is at least
$\max\left\{\frac{k-l}{2k-l}n+\frac{1}{2k-l}\left(\sum_{j=1}^{l-2}j\cdot|M_{j}|\right)+\frac{l-1}{2k-l}\left(\sum_{j=l-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right),~l=1,\ldots,\lfloor\frac{k}{2}\rfloor+1\right\},$
and then for each $l=l^{*}$, it follows that
$\left\{\frac{k-(l^{*}+1)}{2k-(l^{*}+1)}n+\frac{1}{2k-(l^{*}+1)}\left(\sum_{j=1}^{(l^{*}+1)-2}j\cdot|M_{j}|\right)+\frac{(l^{*}+1)-1}{2k-(l^{*}+1)}\left(\sum_{j=(l^{*}+1)-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)\right\}$-
$\left\{\frac{k-l^{*}}{2k-l^{*}}n+\frac{1}{2k-l^{*}}\left(\sum_{j=1}^{l^{*}-2}j\cdot|M_{j}|\right)+\frac{l^{*}-1}{2k-l^{*}}\left(\sum_{j=l^{*}-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)\right\}$
$=$ $\frac{-kn}{[2k-(l^{*}+1)](2k-l^{*})}+\frac{(2k-l^{*})(l^{*}-1)|M_{l^{*}-1}|+\sum_{j=1}^{l^{*}-2}j\cdot|M_{j}|}{[2k-(l^{*}+1)](2k-l^{*})}+$
$\frac{-(2k-l^{*}-1)(l^{*}-1)|M_{l^{*}-1}|+(2k-1)\sum_{j=l^{*}}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|}{[2k-(l^{*}+1)](2k-l^{*})}$
$=$ $\frac{-kn}{[2k-(l^{*}+1)](2k-l^{*})}+\frac{\sum_{j=1}^{l^{*}-1}j\cdot|M_{j}|+(2k-1)\sum_{j=l^{*}}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|}{[2k-(l^{*}+1)](2k-l^{*})},$
which implies that $\frac{k-(l^{*}+1)}{2k-(l^{*}+1)}n+\frac{1}{2k-(l^{*}+1)}\left(\sum_{j=1}^{(l^{*}+1)-2}j\cdot|M_{j}|\right)+\frac{(l^{*}+1)-1}{2k-(l^{*}+1)}\left(\sum_{j=(l^{*}+1)-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)>\frac{k-l^{*}}{2k-l^{*}}n+\frac{1}{2k-l^{*}}\left(\sum_{j=1}^{l^{*}-2}j\cdot|M_{j}|\right)+\frac{l^{*}-1}{2k-l^{*}}\left(\sum_{j=l^{*}-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)$, when $\sum_{j=1}^{l^{*}-1}j\cdot|M_{j}|+(2k-1)(|M_{l^{*}}|+\cdots+|M_{\lfloor\frac{k}{2}\rfloor}|)>kn$. Hence, we can deduce that
  • $\frac{k-2}{2k-2}n+\frac{1}{2k-2}\left(\sum_{j=1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)>\frac{k-1}{2k-1}n$, when $(2k-1)(|M_{1}|+\cdots+|M_{\lfloor\frac{k}{2}\rfloor}|)>kn$,
  • $\frac{k-3}{2k-3}n+\frac{1}{2k-3}|M_{1}|+\frac{2}{2k-3}\left(\sum_{j=2}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)>\frac{k-2}{2k-2}n+\frac{1}{2k-2}\left(\sum_{j=1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)$, when $|M_{1}|+(2k-1)(|M_{2}|+\cdots+|M_{\lfloor\frac{k}{2}\rfloor}|)>kn$,
  • $\ldots$
  • $\frac{k-(\lfloor\frac{k}{2}\rfloor+1)}{2k-(\lfloor\frac{k}{2}\rfloor+1)}n+\frac{1}{2k-(\lfloor\frac{k}{2}\rfloor+1)}\left(\sum_{j=1}^{(\lfloor\frac{k}{2}\rfloor+1)-2}j\cdot|M_{j}|\right)+\frac{(\lfloor\frac{k}{2}\rfloor+1)-1}{2k-(\lfloor\frac{k}{2}\rfloor+1)}\left(\sum_{j=(\lfloor\frac{k}{2}\rfloor+1)-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)>\frac{k-\lfloor\frac{k}{2}\rfloor}{2k-\lfloor\frac{k}{2}\rfloor}n+\frac{1}{2k-\lfloor\frac{k}{2}\rfloor}\left(\sum_{j=1}^{\lfloor\frac{k}{2}\rfloor-2}j\cdot|M_{j}|\right)+\frac{\lfloor\frac{k}{2}\rfloor-1}{2k-\lfloor\frac{k}{2}\rfloor}\left(\sum_{j=\lfloor\frac{k}{2}\rfloor-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)$, when $\sum_{j=1}^{\lfloor\frac{k}{2}\rfloor-1}j\cdot|M_{j}|+(2k-1)|M_{\lfloor\frac{k}{2}\rfloor}|>kn$,
which means that the larger $|M_{j}|,~j=1,\ldots,\lfloor\frac{k}{2}\rfloor$, the larger the general lower bound on the size of control node set, and the influence of the values of $|M_{1}|,\ldots,|M_{\lfloor\frac{k}{2}\rfloor}|$ on the value of the general lower bound is from small to large, that is, $|M_{\lfloor\frac{k}{2}\rfloor}|$ has the greatest impact on the value of the general lower bound.
@@ 6 Boolean networks consisting of nested canalyzing functions
Finally, we consider the case where the nested canalyzing function is assigned to each node, and we refer to this BN as a $k$-$k$-NC-BN. The idea behind the following proposition is to use a kind of shift register again, where the details are substantially different from the previous ones.
@@
When $(n\mod k)=0$, we partition $x_{1},x_{2},\ldots,x_{n}]$ into $\frac{n}{k}$ blocks:
$x_{1},x_{2},\ldots,x_{k}],[x_{k+1},x_{k+2},\ldots,x_{2k}],\ldots,$
@@
Remark 13
On the other hand, there exists a $k$-$k$-NC-BN whose minimum control node set is larger than that given by Theorem 6. For example, we construct a 3-3-NC-BN as follows:
$x_{1}(t+1)=x_{1}(t)\lor(\overline{x_{2}(t)}\land\overline{x_{3}(t)}),$
$x_{2}(t+1)=x_{2}(t)\lor(\overline{x_{3}(t)}\land\overline{x_{4}(t)}),$
$x_{3}(t+1)=x_{3}(t)\lor(\overline{x_{4}(t)}\land\overline{x_{5}(t)}),$
$x_{4}(t+1)=x_{4}(t)\lor(\overline{x_{5}(t)}\land\overline{x_{1}(t)}),$
$x_{5}(t+1)=x_{5}(t)\lor(\overline{x_{1}(t)}\land\overline{x_{2}(t)}).$
Suppose that ${\bf x}^{0}=[1,1,1,1,1]$ and ${\bf x}^{T}=[0,0,0,0,0]$. It is straightforward to see that the size of the minimum control node set must be 55 (i.e., all nodes are control nodes because $x_{i}(t)=1$ would hold for all $t$ regardless of the state of other nodes unless $x_{i}$ is a control node), which is larger than that ($\frac{k-1}{2k-1}n=2$) given by Theorem 6.
@@
7 Simulation results
To evaluate the effectiveness of the derived bounds, we conduct three sets of experiments. The first set includes 100 randomly generated $k$-$k$-XOR-BNs with parameters $(n=10,k=2)$ and $(n=10,k=3)$. The second set includes 100 randomly generated simple $k$-$k$-AND-BNs with parameters $(n=10,k=2)$ and $(n=10,k=3)$. The third set includes 100 randomly generated simple $k$-$k$-NC-BNs with parameters $(n=10,k=2)$ and $(n=10,k=3)$. For each generated BN, we perform a brute-force search to determine the minimum control node set that guarantees controllability. The results, obtained by MATLAB simulations on a MacBook Pro (Apple M4 Pro, 24 GB RAM), are summarized in Fig. 9, which shows that the size of the minimum control node set always lies between the derived lower and upper bounds.
In addition, we find that there exist some $k$-$k$-NC-BNs whose minimum control node sets are smaller than $\frac{n}{k}$, suggesting that the derived best case upper bound may be further improved. Nevertheless, it can still be observed that some $k$-$k$-NC-BNs are easier to control than any simple $k$-$k$-AND-BN.
Figure 9: Comparison between the minimum control node set size and the derived bounds.
@@ 8 Conclusion
  • According to Theorem 2, when $t^{*}=2$, the general lower bound on the size of the control node set is $\frac{n}{2}$. On the other hand, the best case upper bound for 2-2-XOR-BNs is $\frac{n}{2}$ mentioned in Proposition 2, which means that the general lower bound==best case upper bound in this case, i.e., these bounds are optimal.
@@
  • More generally, according to Remark 10, it can be seen that the larger of the values of $|M_{j}|,\;j=1,\ldots,\lfloor\frac{k}{2}\rfloor$, the larger the general lower bound given in Theorem 8, and the influence of the values of $|M_{1}|,\ldots,|M_{\lfloor\frac{k}{2}\rfloor}|$ on the value of the general lower bound is from small to large, that is, $|M_{\lfloor\frac{k}{2}\rfloor}|$ has the greatest impact on the value of the general lower bound.
@@
Notably, all of the above results involving the AND function also apply to the OR function.
However, there are still some open issues that need to be resolved in the future. First, the gap between the bounds given by Proposition 2 and Theorem 3 is significant, so it is necessary to narrow this gap, and some improvements on XOR-BNs [32] were developed based on this work after its submission. Moreover, a general upper bound or a worst case lower bound for $k$-$k$-AND-BN with negation (including the case of $k=2$) is not given in the paper, so it is necessary to derive these bounds further.